The units digit of a two- digit number is one less than 3 times the tens digit. The sum of the digits is 11. Find the number
Answer by mananth(16949)

let ten's digit be x
and unit's digit be y
..
the number = 10x+y
...
y=3x-1
y-3x=-1...................1
x+y =11..................2
multiply equation1 by -1 and add to 2
-y+3x=1
x+y=11
4x= 12
x=3
x+y=11
so y =8
the number is 38
...
m.ananth@hotmail.ca